### v1

# Extrapolation for bilinear compact operators in the variable exponent setting

Abstract.
We establish extrapolation of compactness for bilinear operators in the scale of weighted variable exponent Lebesgue spaces. First, we prove an abstract principle relying on the Cobos–Fernández–Cabrera–Martínez theorem. Then, as an application we deduce new compactness results for the commutators of bilinear $\omega$-Calderón–Zygmund operators and bilinear fractional integrals acting on weighted variable exponent Lebesgue spaces. Our work extends and unifies among others earlier works of the second named author together with Hytönen as well as Oikari.
Key words and phrases:
Weighted extrapolation, multilinear variable weights, compact operators, Calderón–Zygmund operators, fractional integral operators, commutators
2020 Mathematics Subject Classification:
Primary: 42B20, 42B35; Secondary: 46B70, 47H60
S. L. was supported by the Primus research programme PRIMUS/21/SCI/002 of Charles University and the Foundation for Education and European Culture, founded by Nicos and Lydia Tricha.
1. Introduction
This paper brings together several areas: weighted estimates for singular integrals, multilinear operators, variable exponent Lebesgue spaces, and extrapolation of boundedness, respectively compactness. Let us briefly recall the history of each of the involved areas.
It is a classical result that if $T$ is a Calderón–Zygmund operator on $\mathbb{R}^{n}$, $1<p<\infty$, and $w$ is a weight (meaning an a.e. positive function) on $\mathbb{R}^{n}$, such that the so called Muckenhoupt $A_{p}$ characteristic
$w]_{A_{p}}:=\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}w(x)^{p}\,\mathrm{d}x\right)^{1/p}\left(\frac{1}{|Q|}\int_{Q}w(x)^{-p^{\prime}}\,\mathrm{d}x\right)^{1/p^{\prime}}$
is finite, where the supremum is taken over all cubes $Q\subset\mathbb{R}^{n}$ (with faces parallel to the coordinate hyperplanes) and $\frac{1}{p}+\frac{1}{p^{\prime}}=1$, then $T$ acts boundedly from $L^{p}(w)$ into itself. Here we denote
$\|f\|_{L^{p}(w)}:=\left(\int_{\mathbb{R}^{n}}|f(x)w(x)|^{p}\,\mathrm{d}x\right)^{1/p}.$
Such boundedness results and quantitative versions of them are often proved more easily for $p=2$, among other reasons, because $L^{2}(w)$ is a Hilbert space. To extend them to the whole range $1<p<\infty$, one needs an extrapolation result. In this case, the relevant method was introduced by Rubio de Francia [34]. Several works followed further developing this method, until Duoandikoetxea [14] proved a sharp version of it. We refer to the book [9] for a thorough historical overview as well as detailed proofs and generalizations.
There exist several directions for studying more general problems than the above mentioned boundedness. One consists in considering multilinear Calderón–Zygmund operators. Lerner, Ombrosi, Pérez, Torres and Trujillo-González [23] showed that if $T$ is a $m$-linear Calderón–Zygmund operator on $\mathbb{R}^{n}$, $\vec{p}=(p_{1},\ldots,p_{m})$ is a $m$-tuple of exponents from $(1,\infty)$, and $\vec{w}=(w_{1},\ldots,w_{m})$ is a $m$-tuple of weights such that
$\vec{w}]_{A_{\vec{p}}}:=\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}\prod_{j=1}^{m}w_{j}(x)^{p}\,\mathrm{d}x\right)^{1/p}\prod_{j=1}^{m}\left(\frac{1}{|Q|}\int_{Q}w_{j}(x)^{-p_{j}^{\prime}}\,\mathrm{d}x\right)^{1/p_{j}^{\prime}}$
is finite, where
$\frac{1}{p}=\sum_{j=1}^{m}\frac{1}{p_{j}},$
then $T$ maps $L^{p_{1}}(w_{1})\times\ldots\times L^{p_{m}}(w_{m})$ into $L^{p}(\prod_{j=1}^{m}w_{j})$ boundedly. Extrapolation results of such boundedness have played here also a key role in the further development of the theory, see [32].
Another direction consists in considering weighted Lebesgue spaces $L^{p(\cdot)}(w)$ with a variable exponent $p(\cdot)$, that means $p(\cdot)$ is now a not necessarily constant function $p(\cdot):\mathbb{R}^{n}\to[0,\infty]$. For the precise definition of these spaces we refer to Section 2. An exhaustive presentation of the unweighted variable exponent Lebesgue spaces $L^{p(\cdot)}(\mathbb{R}^{n})$ can be found in [12, 13]. Cruz-Uribe, Fiorenza and Neugebauer [8] showed that for every $p(\cdot)\in\mathscr{P}\cap LH$ (see Section 2 for the notation), the Hardy–Littlewood maximal operator is bounded on $L^{p(\cdot)}(w)$ if any only if
$w]_{\mathcal{A}_{p(\cdot)}}:=\sup_{Q}|Q|^{-1}\|w\chi_{Q}\|_{p(\cdot)}\|w^{-1}\chi_{Q}\|_{p^{\prime}(\cdot)}$
is finite. This was extended in [33] to the bilinear setting, where moreover analogous boundedness results for bilinear singular integrals were proved. Also, there is by now a rich extrapolation theory for variable exponent spaces both in the linear and the multilinear setting, see [11, 4, 31]. However, it is still unknown if one can prove the exact linear and multilinear analogue of Rubio de Francia’s extrapolation of boundedness on weighted Lebesgue spaces with variable exponents (see also [22, Remark 1.7]).
Over the last few years, the extension of extrapolation results to compact operators has drawn the attention of many authors. In particular, the first work that dealt with this mathematical problem in the linear setting was due to Hytönen and the second named author [17]. Further extensions of this result in the setting of weighted Morrey spaces, two weight problems, Banach function spaces and one sided situations can be found in [21], [24], [26] (see also [22] for an alternative proof in the context of weighted Lebesgue spaces with variable exponents) and [29], respectively. Moreover, first Cao, Olivo and Yabuta [5] established extrapolation results for multilinear compact operators (see also [18]).
Main results
Motivated by the previously mentioned works, in this paper we establish the following abstract diagonal and off-diagonal extrapolation theorem for bilinear compact operators in the setting of weighted variable exponent Lebesgue spaces (see Section 2 for detailed definitions).
1.1 Theorem.
Let $T$ be a bilinear operator such that there exist constants $t\in[1,\infty)$ and $\gamma\in\left[0,\frac{2}{t}\right)$ satisfying both of the following two assumptions.
  1. (1)
    For all $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, $q_{0}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ with
    $t<\min\{(p_{0,j})_{-}:~j=1,2\}$
    and
    $\frac{1}{p_{0}(\cdot)}=\frac{1}{p_{0,1}(\cdot)}+\frac{1}{p_{0,2}(\cdot)},\quad\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$
    the following holds. Whenever $\vec{w}_{0}$ is a $2$-tuple of weights such that $\vec{w}_{0}\in\mathcal{A}^{\gamma}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}$ and $\vec{w}_{0}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}_{0}(\cdot)}{t},\frac{q_{0}(\cdot)}{t}}$, we have
    $T:L^{p_{0,1}(\cdot)}(w_{0,1})\times L^{p_{0,2}(\cdot)}(w_{0,2})\to L^{q_{0}(\cdot)}(\nu_{\vec{w}_{0}})$
    boundedly.
  2. (2)
    There exist some fixed $2$-tuple of variable exponents $\vec{p}_{1}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, a fixed variable exponent $q_{1}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ with
    $t<\min\{(p_{1,j})_{-}:~j=1,2\}$
    and
    $\frac{1}{p_{1}(\cdot)}=\frac{1}{p_{1,1}(\cdot)}+\frac{1}{p_{1,2}(\cdot)},\quad\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\gamma,$
    and a fixed $2$-tuple of weights $\vec{w}_{1}$ with $\vec{w}_{1}\in\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}$ and $\vec{w}_{1}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}_{1}(\cdot)}{t},\frac{q_{1}(\cdot)}{t}}$, such that
    $T:L^{p_{1,1}(\cdot)}(w_{1,1})\times L^{p_{1,2}(\cdot)}(w_{1,2})\to L^{q_{1}(\cdot)}(\nu_{\vec{w}_{1}})$
    compactly.
Then: For all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, $q(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$, such that
$t<\min\{(p_{j})_{-}:~j=1,2\}$
and
$\frac{1}{p(\cdot)}=\frac{1}{p_{1}(\cdot)}+\frac{1}{p_{2}(\cdot)},\quad\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma,$
and for all $2$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$, we have
$T:L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})\to L^{q(\cdot)}(\nu_{\vec{w}})$
compactly.
As applications, we obtain new compactness results for the commutators of bilinear $\omega$-Calderón–Zygmund operators and bilinear fractional integrals.
1.2 Remark.
By assuming $t=1$ and choosing all the variable exponents to be constant ones in the assumptions and conclusion of Theorem 1.1, we can extend or recover several results of [5, 18, 22, 26].
The remainder of this paper is organized as follows. In Section 2, we recall the definition of bilinear compact operators and state several definitions about unweighted and weighted variable Lebesgue spaces. In Section 3, we gather some auxiliary lemmas that will aid us in proving our main result. It is worth mentioning that the proof of one of these lemmas, namely, the direction “if” of Lemma 3.4 is new. Section 4 is devoted to the proof of two key factorization results in the setting of variable multilinear Muckenhoupt weights (see Lemmas 4.1 and 4.2). In Section 5, we recall several known facts about complex interpolation for bilinear compact operators. Among them is the abstract result of Cobos–Fernández-Cabrera–Martínez [7, Theorem 3.2]. Lastly, in Sections 6 and 7, we provide the proof of Theorem 1.1 and its applications, respectively.
Notation
We write $x\lesssim y$ to mean that there exists a constant $C>0$ such that $x\leq Cy$. This constant $C$ is just an absolute constant or one depending on parameters that are specified each time or that are understood by the context. Moreover, we denote by $|Q|$ the n-dimensional Lebesgue measure of a cube $Q\subset\mathbb{R}^{n}$ and $\langle w\rangle_{Q}$ will denote the average of $w$ over $Q$, that is, $|Q|^{-1}\int_{Q}w(x)\,\mathrm{d}x$.
2. Preliminaries
Throughout the paper, we mostly consider bilinear compact operators. Hence, in this section we recall this definition for the reader’s convenience. Given the normed spaces $X,Y$ and $Z$, a bilinear operator $T:X\times Y\rightarrow Z$ is said to be compact if the set $\{T(x,y):\;\|x\|\leq 1,\;\;\|y\|\leq 1\}$ is precompact in $Z$. Some other equivalent formulations of this definition were given in [2, Proposition 1].
2.A. Variable Lebesgue spaces
In this subsection, we recall some basic definitions of variable Lebesgue spaces. For more information, we refer the reader to [12] and [13].
2.A.1. Notation for variable exponent functions
Given any measurable function $p(\cdot):\mathbb{R}^{n}\to[0,\infty]$ and a set $E\subset\mathbb{R}^{n}$, we denote
$p_{-}(E)=\operatorname*{ess\;inf}_{x\in E}p(x),\qquad p_{+}(E)=\operatorname*{ess\;sup}_{x\in E}p(x),$
and if $E=\mathbb{R}^{n}$, then we will write $p_{-}(\mathbb{R}^{n})=p_{-}$ and $p_{+}(\mathbb{R}^{n})=p_{+}$.
In this paper, we will consider several special classes of measurable functions $p(\cdot):\mathbb{R}^{n}\to[0,\infty]$:
$\mathcal{P}_{0}:=\big\{\text{all Lebesgue measurable functions }p(\cdot):\mathbb{R}^{n}\to(0,\infty]\big\},$
$\mathscr{P}_{0}:=\{p(\cdot)\in\mathcal{P}_{0}:~p_{+}<\infty\},$
$\mathcal{P}:=\{p(\cdot)\in\mathcal{P}_{0}:~p(x)\geq 1,\text{ for all }x\in\mathbb{R}^{n}\},$
and lastly
$\mathscr{P}:=\{p(\cdot)\in\mathcal{P}:~1<p_{-}\leq p_{+}<\infty\}.$
Given $p(\cdot)\in\mathcal{P}$, the dual exponent $p^{\prime}(\cdot)\in\mathcal{P}$ is given by $1/p(\cdot)+1/p^{\prime}(\cdot)=1.$
If $0<|E|<\infty$, we define the harmonic mean $p_{E}$ on the set $E\subset\mathbb{R}^{n}$ by
[…]
2.A.2. Luxemburg norm of variable Lebesgue spaces
Given $p(\cdot)\in\mathcal{P}_{0}$, define the modular associated with $p(\cdot)$ by
(2.1) $\rho_{p(\cdot)}(f):=\int_{\mathbb{R}^{n}\setminus\Omega_{\infty}}|f(x)|^{p(x)}\,\mathrm{d}x+\|f\|_{L^{\infty}(\Omega_{\infty})},$
where $\Omega_{\infty}:=\{x\in\mathbb{R}^{n}\;:\;p(x)=\infty\}$. The variable Lebesgue space $L^{p(\cdot)}(\mathbb{R}^{n})$ consists of all measurable functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that
$\|f\|_{L^{p(\cdot)}}=\|f\|_{{p(\cdot)}}:=\inf\{\lambda\in(0,\infty):\rho_{p(\cdot)}(f/\lambda)\leq 1\}<\infty.$
In what follows, we always assume that $p_{+}<\infty$ and hence the latter term in (2.1) is zero.
2.A.3. Weighted variable Lebesgue spaces
By a weight $w$ we mean a measurable function $w$ on $\mathbb{R}^{n}$ such that $0<w(x)<\infty$ for almost every $x\in\mathbb{R}^{n}$. Given a weight $w$, we define the weighted variable Lebesgue space $L^{p(\cdot)}(w)$ through
$\|f\|_{L^{p(\cdot)}(w)}:=\|fw\|_{p(\cdot)}<\infty.$
2.A.4. Log-Hölder continuity
For technical reasons, we will be considering variable exponent functions possessing the local and asymptotic Hölder continuity properties. We recall these below.
2.2 Definition.
For $p(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$, we say that
  1. (1)
    it is locally log-Hölder continuous, denoted by $p(\cdot)\in LH_{0},$ if there exists $C_{0}>0$ such that for all $x,y\in\mathbb{R}^{n}$, whenever $|x-y|<\tfrac{1}{2}$, then
    $|p(x)-p(y)|\leq\frac{C_{0}}{-\log(|x-y|)};$
  2. (2)
    it is log-Hölder continuous at infinity, denoted by $p(\cdot)\in LH_{\infty},$ if there exist $p_{\infty}\in\mathbb{R}$ and $C_{\infty}>0$ such that for all $x\in\mathbb{R}^{n}$,
    $|p(x)-p_{\infty}|\leq\frac{C_{\infty}}{\log(e+|x|)}.$
We define $LH:=LH_{0}\cap LH_{\infty}.$
2.B. Variable exponent Muckenhoupt characteristics
In what follows, we work directly in the off-diagonal setting, namely, we consider $\gamma\in(0,1)$ and not only $\gamma=0$. Let $\gamma\in[0,1)$ be fixed and denote
(2.3) $\mathcal{E}_{(1,\infty)}^{0}$ $:=LH\cap\mathscr{P},$
$\mathcal{E}^{\gamma}_{(1,\infty)}$ $:=\left\{\big(p(\cdot),q(\cdot)\big)\in\mathcal{E}_{(1,\infty)}^{0}\times\mathcal{E}_{(1,\infty)}^{0}:1/p(\cdot)-1/q(\cdot)=\gamma\right\}.$
The appropriate class of weights is as follows.
2.4 Definition ([11], Definition 2.10).
Let $\gamma\in[0,1)$ and $p(\cdot),q(\cdot)\in\mathcal{P}$ be such that $1/p(\cdot)-1/q(\cdot)=\gamma.$ Then, a weight $w$ belongs to the class $\mathcal{A}_{p(\cdot),q(\cdot)}^{\gamma},$ if
$w]_{\mathcal{A}_{p(\cdot),q(\cdot)}^{\gamma}}:=\sup_{Q}|Q|^{\gamma-1}\|w\chi_{Q}\|_{q(\cdot)}\|w^{-1}\chi_{Q}\|_{p^{\prime}(\cdot)}<\infty,$
where the supremum is taken over all cubes.
When $\gamma=0$ we understand that $p(\cdot)\in\mathcal{E}_{(1,\infty)}^{0}$ is equivalent to $(p(\cdot),p(\cdot))\in\mathcal{E}_{(1,\infty)}^{0}$ and $\mathcal{A}_{p(\cdot),q(\cdot)}^{0}=\mathcal{A}_{p(\cdot)}$.
2.5 Definition.
Given variable exponents $p_{1}(\cdot),\dots,p_{m}(\cdot)$, $p_{0,1}(\cdot),\dots,p_{0,m}(\cdot)$ and $p_{1,1}(\cdot),\dots,\\ p_{1,m}(\cdot)$ we define the vectors
$\vec{p}(\cdot):=(p_{1}(\cdot),\dots,p_{m}(\cdot)),\quad\vec{p}_{0}(\cdot):=(p_{0,1}(\cdot),\dots,p_{0,m}(\cdot))\quad\text{and}\quad\vec{p}_{1}(\cdot):=(p_{1,1}(\cdot),\dots,p_{1,m}(\cdot)).$
2.6 Definition.
Given vectors of weights $\vec{w}=(w_{1},\dots,w_{m})$, $\vec{w}_{0}=(w_{0,1},\dots,w_{0,m})$ and $\vec{w}_{1}=(w_{1,1},\dots,w_{1,m})$ we define
$\nu_{\vec{w}}:=\prod_{j=1}^{m}w_{j},\qquad\nu_{\vec{w}_{0}}:=\prod_{j=1}^{m}w_{0,j}\qquad\text{and}\qquad\nu_{\vec{w}_{1}}:=\prod_{j=1}^{m}w_{1,j}.$
The following class of fractional-type variable multiple weights was first introduced in [6].
2.7 Definition ([6], Definition 1.3).
Let $\gamma\in[0,m)$ be a constant. Let $p_{j}(\cdot)\in\mathscr{P}$, $j=1,\dots,m$ and let $q\in\mathscr{P}_{0}$ such that $\gamma=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}$, where we define $p\in\mathscr{P}_{0}$ by $\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}$. We say that a vector of weights $\vec{w}=(w_{1},\dots,w_{m})$ satisfies the multilinear $\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ condition (or $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$) if
$\vec{w}]_{\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}}:=\sup_{Q}|Q|^{\gamma-m}\|\nu_{\vec{w}}\chi_{Q}\|_{q(\cdot)}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{p_{j}^{\prime}(\cdot)}<\infty,$
where the supremum is taken over all cubes $Q\subset\mathbb{R}^{n}$.
2.8 Remark.
If $\gamma=0$, then $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}^{\gamma}=\mathcal{A}_{\vec{p}(\cdot)}$. This bi/multilinear version of the variable $\mathcal{A}_{\vec{p}(\cdot)}$ condition was first introduced in [33]. Moreover, notice that in the above definitions we can substitute cubes for balls.
3. Auxiliary results
In this section, we recall several known results from the theory of variable Lebesgue spaces that we will need for our purposes.
3.1 Lemma ([12], Proposition 2.18).
Given a measurable function $p(\cdot):\mathbb{R}^{n}\rightarrow(0,\infty)$ such that $p_{+}<\infty$, there holds that $\||f|^{s}\|_{p(\cdot)}=\|f\|^{s}_{sp(\cdot)},$ for all $s>0.$
The following Hölder’s inequality hold in variable Lebesgue spaces.
3.2 Lemma ([12], Corollary 2.30).
Fix a positive integer $m\geq 2$ and let $p_{1}(\cdot),\dots,p_{m}(\cdot)\in\mathcal{P}$ satisfy
$\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}=1,$
where $j=1,\dots,m$. Then, for all $f_{j}\in L^{p_{j}(\cdot)}$, one has
$\int_{\mathbb{R}^{n}}|f_{1}\cdots f_{m}|\lesssim\|f_{1}\|_{L^{p_{1}(\cdot)}}\cdots\|f_{m}\|_{L^{p_{m}(\cdot)}},$
where the implicit constant depends only on the $p_{j}(\cdot)$ and $m.$
We recall the following corollary of [10, Theorem 1.1] as stated in [22, Lemma 2.8].
3.3 Lemma ([10], Corollary 4.3).
Let $r(\cdot)\in LH\cap\mathcal{P}$ with $r_{+}<\infty$. If $u\in\mathcal{A}_{r(\cdot)}$, then there exists an exponent $s>1$ depending only on $n,$ $r(\cdot)$ and $u]_{\mathcal{A}_{r(\cdot)}}$ such that for all $t\in(1,s)$ and for all cubes cubes $Q\subset\mathbb{R}^{n}$,
$|Q|^{-\frac{1}{{tr_{Q}}}}\|u\chi_{Q}\|_{L^{tr(\cdot)}}\lesssim|Q|^{-\frac{1}{r_{Q}}}\|u\chi_{Q}\|_{L^{r(\cdot)}},$
where the implicit constant depends only on $n$, $r(\cdot)$, $u]_{\mathcal{A}_{r(\cdot)}}$ and $1]_{\mathcal{A}_{\widetilde{v}(\cdot)}}$$\infty$, where
$\widetilde{v}(\cdot):=\frac{r(\cdot)}{\frac{1}{t}-\frac{1}{s}}.$
Now, we emphasize that it is possible to characterize the multilinear $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ weights in terms of the $\mathcal{A}_{p(\cdot)}$ condition. In particular, relying on arguments as in [6, Lemma 1.2] and [33, Proposition 4.7] we obtain the following result.
3.4 Lemma.
Let $p_{j}\in\mathscr{P}$, $j=1,\dots,m,$ with $\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}$ and $\gamma=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}\in[0,m)$. Then $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ if and only if
(3.5) $\begin{cases}w_{j}^{-\frac{1}{m}}\in\mathcal{A}_{mp_{j}^{\prime}(\cdot)},&j=1,\dots,m,\\ \nu_{\vec{w}}^{\frac{1}{m}}\in\mathcal{A}_{mq(\cdot)}.\end{cases}$
Proof.
Notice that the proof of “only if” direction is contained in [6, Lemma 1.2]. To prove the “if” direction, we assume that (3.5) holds. Then, by Lemma 3.1 and Hölder’s inequality (Lemma 3.2) with exponents
$\frac{1}{\frac{\gamma+m^{2}}{m}(mq)^{\prime}(\cdot)}+\sum_{j=1}^{m}\frac{1}{\frac{\gamma+m^{2}}{m}(mp_{j}^{\prime})^{\prime}(\cdot)}=1$
we get
[…]
$\lesssim|Q|^{-\frac{\gamma}{m}-m}\|(\nu_{\vec{w}}^{-\frac{1}{m}})^{\frac{m}{\gamma+m^{2}}}\chi_{Q}\|_{\frac{\gamma+m^{2}}{m}(mq)^{\prime}(\cdot)}^{\frac{\gamma+m^{2}}{m}}\prod_{j=1}^{m}\|(w_{j}^{\frac{1}{m}})^{\frac{m}{\gamma+m^{2}}}\chi_{Q}\|_{\frac{\gamma+m^{2}}{m}(mp_{j}^{\prime})^{\prime}(\cdot)}^{\frac{\gamma+m^{2}}{m}}$
$=|Q|^{-\frac{\gamma}{m}-m}\|\nu_{\vec{w}}^{-\frac{1}{m}}\chi_{Q}\|_{(mq)^{\prime}(\cdot)}\prod_{j=1}^{m}\|w_{j}^{\frac{1}{m}}\chi_{Q}\|_{(mp_{j}^{\prime})^{\prime}(\cdot)}.$
Hence, this and Definition 2.4 of $\mathcal{A}_{p(\cdot),q(\cdot)}^{0}=\mathcal{A}_{p(\cdot)}$ imply that
$\|\nu_{\vec{w}}\chi_{Q}\|_{q(\cdot)}^{\frac{1}{m}}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{p_{j}^{\prime}(\cdot)}^{\frac{1}{m}}$ $=\|\nu_{\vec{w}}^{\frac{1}{m}}\chi_{Q}\|_{mq(\cdot)}\prod_{j=1}^{m}\|w_{j}^{-\frac{1}{m}}\chi_{Q}\|_{mp_{j}^{\prime}(\cdot)}$
$\lesssim|Q|^{-\frac{\gamma}{m}-m}\|\nu_{\vec{w}}^{\frac{1}{m}}\chi_{Q}\|_{mq(\cdot)}\|\nu_{\vec{w}}^{-\frac{1}{m}}\chi_{Q}\|_{(mq)^{\prime}(\cdot)}$
$\qquad\times\prod_{j=1}^{m}\|w_{j}^{\frac{1}{m}}\chi_{Q}\|_{(mp_{j}^{\prime})^{\prime}(\cdot)}\|w_{j}^{-\frac{1}{m}}\chi_{Q}\|_{mp_{j}^{\prime}(\cdot)}$
$\lesssim|Q|^{-\frac{\gamma}{m}-m}|Q||Q|^{m}.$
Thus,
$\bigg(|Q|^{\gamma-m}\|\nu_{\vec{w}}\chi_{Q}\|_{q(\cdot)}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{p_{j}^{\prime}(\cdot)}\bigg)^{\frac{1}{m}}\lesssim 1.$
This and Definition 2.7 of $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ allow us to conclude that $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$, which completes the proof. ∎
3.6 Remark.
If $m=2$ and $\gamma=0$, then Lemma 3.4 is exactly [33, Proposition 4.7].
4. Factorization results for multilinear Muckenhoupt weights with variable exponents
One of the key ingredients for the proof of Theorem 1.1 is the following factorization result for the variable multilinear Muckenhoupt weights. The proof of this result is based on similar ideas of [18].
4.1 Lemma.
Let $\gamma\in[0,m)$ be a fixed constant. Let $\vec{p}_{1}(\cdot),\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$, $q_{1}(\cdot),q(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}$ with
$\frac{1}{p_{1}(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{1,j}(\cdot)},\quad\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)},\quad\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma.$
Let also $\vec{w}_{1}\in\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}$ and $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$. Given any $\theta\in(0,1)$, define $\vec{p}_{0}(\cdot)\in(\mathcal{P}_{0})^{m}$, $q_{0}(\cdot)\in\mathcal{P}_{0}$ by
$\frac{1}{p_{j}(\cdot)}$ $=\frac{1-\theta}{p_{0,j}(\cdot)}+\frac{\theta}{p_{1,j}(\cdot)},\quad j=1,\ldots,m,$
$\frac{1}{q(\cdot)}$ $=\frac{1-\theta}{q_{0}(\cdot)}+\frac{\theta}{q_{1}(\cdot)},$
as well as the vector of weights $\vec{w}_{0}$ by
$w_{j}$ $=w_{0,j}^{1-\theta}w_{1,j}^{\theta},\quad j=1,\ldots,m.$
Then, there exist $\eta>0$, depending only on $n,m,$ $\vec{p}_{1}(\cdot),\vec{p}(\cdot),q_{1}(\cdot),q(\cdot)$ and the Muckenhoupt characteristics $\vec{w}_{1}]_{\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}}$, $\vec{w}]_{\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}},$ such that for all $\theta\in(0,\eta)$ all of the following hold:
  1. (1)
    $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$ $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0},$ and $\frac{1}{p_{0}(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{0,j}(\cdot)}$, $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma$.
  2. (2)
    If $(q_{1})_{-},q_{-}>1$, then also $(q_{0})_{-}>1.$
  3. (3)
    $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}^{\gamma}.$
Proof.
Notice that once $\theta$ is fixed, $\vec{p}_{0}(\cdot),q_{0}(\cdot)$ and $\vec{w}_{0}(\cdot)$ are uniquely determined.
By the first part of the proof of [22, Lemma 3.1] we have that for all $j=1,\ldots,m$, by choosing
$\eta<\min\left(\frac{(p_{1,j})_{-}}{(p_{j})_{+}},1-\frac{1}{(p_{j})_{-}}\right),$
we can ensure that $p_{0,j}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}.$
Furthermore, it is easy to see that $mq(\cdot),mq_{1}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ and
$\frac{1}{mq(\cdot)}=\frac{1-\theta}{mq_{0}(\cdot)}+\frac{\theta}{mq_{1}(\cdot)},$
with $mq_{0}(\cdot)$ being completely determined from $\theta$, $mq_{1}(\cdot)$ and $mq(\cdot)$. Thus, by the same token, by choosing $\eta$ small enough depending only on $m$, $q_{1}(\cdot)$ and $q(\cdot)$ we can ensure that $mq_{0}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$, so in particular $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}.$
To sum up, by choosing $\eta$ small enough depending only on $\vec{p}_{1}(\cdot),\vec{p}(\cdot),q_{1}(\cdot),q(\cdot)$ we can ensure that $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ and $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}.$ If by chance $(q_{1})_{-},q_{-}>1,$ then by picking $\eta$ even smaller we can ensure in addition $(q_{0})_{-}>1.$
It is trivial to verify the relation $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma.$
To prove that for small enough $\eta$, $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}^{\gamma}$, it suffices by Lemma 3.4 to show that
$\begin{cases}w_{0,j}^{-\frac{1}{m}}\in\mathcal{A}_{mp_{0,j}^{\prime}(\cdot)},\quad j=1,\ldots,m\\ \nu_{\vec{w}_{0}}^{\frac{1}{m}}\in\mathcal{A}_{mq_{0}(\cdot)}.\end{cases}$
Since $\vec{w}_{1}\in\mathcal{A}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}^{\gamma}$ and $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}^{\gamma}$, by Lemma 3.4 we have
$\begin{cases}w_{1,j}^{-\frac{1}{m}}\in\mathcal{A}_{mp_{1,j}^{\prime}(\cdot)},\quad j=1,\ldots,m\\ \nu_{\vec{w}_{1}}^{\frac{1}{m}}\in\mathcal{A}_{mq_{1}(\cdot)}.\end{cases}$
and
$\begin{cases}w_{j}^{-\frac{1}{m}}\in\mathcal{A}_{mp_{j}^{\prime}(\cdot)},\quad j=1,\ldots,m\\ \nu_{\vec{w}}^{\frac{1}{m}}\in\mathcal{A}_{mq(\cdot)}.\end{cases}$
Moreover, by the proof of [6, Lemma 1.2] we have that the associated variable exponent Muckenhoupt characteristics are all bounded from above by $\vec{w}_{1}]_{\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}}^{\frac{1}{m}},$ respectively $\vec{w}]_{\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}}^{\frac{1}{m}},$ up to implied constants depending only on $n,m,$ as well as $\vec{p}_{1}(\cdot)$ and $q_{1}(\cdot)$, respectively $\vec{p}(\cdot)$ and $q(\cdot).$
Fix now $j=1,\ldots,m.$ Set $v:=w_{j}^{-\frac{1}{m}}$ and $r(\cdot):=mp_{j}^{\prime}(\cdot),$ extending these notations analogously to their endpoint counterparts. Observe that $r(\cdot),r_{1}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$, $v_{1}\in\mathcal{A}_{r_{1}(\cdot)},$ $v\in\mathcal{A}_{r(\cdot)},$
$v=w_{j}^{-\frac{1}{m}}=(w_{0,j}^{-\frac{1}{m}})^{1-\theta}(w_{1,j}^{-\frac{1}{m}})^{\theta}=v_{0}^{1-\theta}v_{1}^{\theta},$
and
$\frac{1}{r(\cdot)}$ $=\frac{1}{mp_{j}(\cdot)^{\prime}}=\frac{1}{m}\left(1-\frac{1}{p_{j}(\cdot)}\right)=\frac{1}{m}\left(1-\theta+\theta-\frac{1-\theta}{p_{0,j}(\cdot)}-\frac{\theta}{p_{1,j}(\cdot)}\right)$
$=\frac{1-\theta}{mp_{0,j}^{\prime}(\cdot)}+\frac{\theta}{mp_{1,j}^{\prime}(\cdot)}=\frac{1-\theta}{r_{0}(\cdot)}+\frac{\theta}{r_{1}(\cdot)}.$
Thus, a direct application of [22, Lemma 3.1] in the diagonal setting and the named variable exponents and weights shows that we can pick $\eta$ small enough, depending only on $n,$ $r_{1}(\cdot),r(\cdot),$ $v_{1}]_{\mathcal{A}_{r_{1}(\cdot)}}$ and $v]_{\mathcal{A}_{r(\cdot)}}$, so that for any $\theta<\eta$ one has $v_{0}\in\mathcal{A}_{r_{0}(\cdot)}.$ Note that here we only need a somewhat simplified version of the second part of the proof of [22, Lemma 3.1]: by considering only the diagonal case, one avoids having to use [11, Proposition 5.4], so that a direct application of the reverse Hölder inequality, Lemma 3.3 is possible.
Similarly, one can ensure that $\nu_{\vec{w}_{0}}^{\frac{1}{m}}\in\mathcal{A}_{mq_{0}(\cdot)}.$ This concludes the proof. ∎
In view of later applications of our extrapolation results, we also need a slightly varied version of Lemma 4.1. Even though the changes are restricted essentially to the notational level, we include a precise statement and a short explanation for the sake of clarity.
4.2 Lemma.
Let $t$ be a fixed constant with
$1<t<\min\left(\min\{(p_{1,j})_{-}:~j=1,\ldots,m\},\min\{(p_{j})_{-}:~j=1,\ldots,m\}\right),$
and let $\gamma\in\left[0,\frac{m}{t}\right)$ be a fixed constant. Consider $\vec{p}_{1}(\cdot),\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ and $q_{1}(\cdot),q(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}$ with
$\frac{1}{p_{1}(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{1,j}(\cdot)},\quad\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)},\quad\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma.$
Let also $\vec{w}_{1}\in\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}$ and $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ such that $\vec{w}_{1}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}_{1}(\cdot)}{t},\frac{q_{1}(\cdot)}{t}}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$.
Given any $\theta\in(0,1)$, define $\vec{p}_{0}(\cdot)\in(\mathcal{P}_{0})^{m}$, $q_{0}(\cdot)\in\mathcal{P}_{0}$ by
$\frac{1}{p_{j}(\cdot)}$ $=\frac{1-\theta}{p_{0,j}(\cdot)}+\frac{\theta}{p_{1,j}(\cdot)},\quad j=1,\ldots,m,$
$\frac{1}{q(\cdot)}$ $=\frac{1-\theta}{q_{0}(\cdot)}+\frac{\theta}{q_{1}(\cdot)},$
as well as the vector of weights $\vec{w}_{0}$ by
$w_{j}$ $=w_{0,j}^{1-\theta}w_{1,j}^{\theta},\quad j=1,\ldots,m.$
Then, there exists $\eta>0$, depending only on $n,m,$ $\vec{p}_{1}(\cdot),\vec{p}(\cdot),q_{1}(\cdot),q(\cdot),$ $t$ and the Muckenhoupt characteristics $\vec{w}_{1}]_{\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}}$, $\vec{w}]_{\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}},$ such that for all $\theta\in(0,\eta)$ all of the following hold:
  1. (1)
    $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$ $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0},$ and $\frac{1}{p_{0}(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{0,j}(\cdot)}$, $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma.$
  2. (2)
    If $(q_{1})_{-},q_{-}>1$, then also $(q_{0})_{-}>1$.
  3. (3)
    $t<\min\{(p_{0,j})_{-}:~j=1,\ldots,m\}.$
  4. (4)
    $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}^{\gamma}.$
  5. (5)
    $\vec{w}_{0}^{t}\in\mathcal{A}_{\frac{\vec{p}_{0}(\cdot)}{t},\frac{q_{0}(\cdot)}{t}}^{t\gamma}$.
Proof.
Set $\delta:=t\gamma\in[0,m)$ and define the $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts
$\vec{\omega}:=\vec{w}^{t},$
$\vec{\pi}(\cdot):=\frac{\vec{p}(\cdot)}{t},$
$\kappa(\cdot):=\frac{q(\cdot)}{t},$
and extend these notation analogously to their counterparts for the endpoint 1. Then, we clearly have
$\vec{\pi}_{1}(\cdot),\vec{\pi}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$
$\kappa_{1}(\cdot),\kappa(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}$
$\vec{\omega}_{1}\in\mathcal{A}^{\delta}_{\vec{\pi}_{1}(\cdot),\kappa_{1}(\cdot)},$
$\vec{\omega}\in\mathcal{A}^{\delta}_{\vec{\pi}(\cdot),\kappa(\cdot)}.$
Moreover, it holds
$\frac{1}{\pi_{1}(\cdot)}-\frac{1}{\kappa_{1}(\cdot)}=\frac{1}{\ \pi(\cdot)}-\frac{1}{\kappa(\cdot)}=\delta.$
The $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts for the endpoint 0 are defined directly in terms of the already defined $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” coounterparts by
$\frac{1}{\pi_{j}(\cdot)}$ $=\frac{1-\theta}{\pi_{0,j}(\cdot)}+\frac{\theta}{\pi_{1,j}(\cdot)},\quad j=1,\ldots,m,$
$\frac{1}{\kappa(\cdot)}$ $=\frac{1-\theta}{\kappa_{0}(\cdot)}+\frac{\theta}{\kappa_{1}(\cdot)},$
$\omega_{j}$ $=\omega_{0,j}^{1-\theta}\omega_{1,j}^{\theta},\quad j=1,\ldots,m.$
Thus, by applying Lemma 4.1 both for the original objects as well as their $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts, we can achieve that by picking $\eta$ small enough, depending on the same quantities as in Lemma 4.1 and in addition $t$, the relation $\theta<\eta$ already ensures that
  • $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$ $q_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0},$ $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$ $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}^{\gamma}$, and in addition $(q_{1})_{-},q_{-}>1\Rightarrow(q_{0})_{-}>1,$
  • $\vec{\pi}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m},$ $\kappa_{0}(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0},$ $\frac{1}{\pi_{0}(\cdot)}-\frac{1}{\kappa_{0}(\cdot)}=\delta,$ $\vec{\omega}_{0}\in\mathcal{A}_{\vec{\pi}_{0}(\cdot),\kappa_{0}(\cdot)}^{\delta}$.
Therefore, it now remains to check that our definition of the $\pi\text{-}\kappa\text{-}\delta\text{-}\omega$” counterparts for the endpoint 0 automatically yields the desired $t$-scaling, that is that we indeed have
(4.3) $\vec{\pi}_{0}(\cdot)=\frac{\vec{p}_{0}(\cdot)}{t},\quad\kappa_{0}(\cdot)=\frac{q(\cdot)}{t},\quad\vec{\omega}_{0}=\vec{w}_{0}^{t}.$
Moreover, we need to check that we automatically have
(4.4) $t<\min\{(p_{0,j})_{-}:~j=1,\ldots,m\}.$
To check (4.3), observe that for all $j=1,\ldots,m$ we have
$\frac{1}{\pi_{0,j}(\cdot)}=\frac{1}{1-\theta}\left(\frac{1}{\pi_{j}(\cdot)}-\frac{\theta}{\pi_{1,j}(\cdot)}\right)=\frac{1}{1-\theta}\left(\frac{t}{p_{j}(\cdot)}-\frac{t\theta}{p_{1,j}(\cdot)}\right)=\frac{t}{p_{0,j}(\cdot)}.$
The other parts of (4.3) are similarly verified. Lastly, for checking (4.4), we have
$(\pi_{0,j})_{-}>1,\quad\forall j=1,\ldots,m$
therefore
$\left(\frac{p_{0,j}}{t}\right)_{-}>1,\quad\forall j=1,\ldots,m,$
in other words
$(p_{0,j})_{-}>t,\quad\forall j=1,\ldots,m,$
concluding the proof. ∎
5. Lions–Peetre compactness theorem in the bilinear setting
In this section we review the basic abstract tool about interpolation for bilinear compact operators that will be used in the proof of Theorem 1.1. For the reader’s convenience, we include some background. None of the material here is new, but we carefully track the assumptions of the various definitions.
5.A. Banach function spaces
We begin with briefly recalling the notion of Banach function spaces. We follow the presentation and definitions from the recent paper [25] of Lorist–Nieraeth, where we refer to for more details, history and proofs.
5.1 Definition.
Let $(\Omega,\mu)$ be a $\sigma$-finite measure space. We denote by $L^{0}(\Omega)$ the space of all measurable functions on $\Omega$. Let $X$ be a Banach space whose underlying set is a subset of $L^{0}(\Omega)$. Following [25], the space $X$ is said to be a Banach function space over $(\Omega,\mu)$ if it satisfies both of the following two properties:
  • Ideal property: If $f\in X$ and $g\in L^{0}(\Omega)$ with $|g|\leq|f|$ $\mu$-a.e., then $g\in X$ and $\|g\|_{X}\leq\|f\|_{X}$.
  • Saturation property: For every measurable set $E\subseteq\Omega$ with $\mu(E)>0$, there exists a measurable set $F\subseteq E$ with $\mu(F)>0$ and $\chi_{F}\in X$.
5.2 Remark.
In several places of the literature, the definition of Banach function spaces includes a much stronger property than the saturation property, namely, that the characteristic function of any measurable set of finite measure belongs to the space. However, as explained in [25], this property is too restrictive, as it excludes in general for example weighted Lebesgue spaces, even with constant exponent.
Several important Banach functions spaces enjoy in addition the Fatou property. While sometimes in the literature it furnished part of the definition of Banach functions spaces, for us, following [25], it will be an additional, desirable property. We follow the formulation in [25].
5.3 Definition.
Let $X$ be a Banach function space over a $\sigma$-finite measure space $(\Omega,\mu)$. We say that $X$ has the Fatou property, if for any pointwise increasing sequence $(f_{k})_{k=1}^{\infty}$ of nonnegative functions in $X$ with $f_{k}\to f$ pointwise $\mu$-a.e. as $k\to\infty$ for some $f\in L^{0}(\Omega)$ and $\sup\limits_{k=1,2,\ldots}\|f_{k}\|_{X}<\infty$, it holds $f\in X$ and $\|f\|_{X}=\lim\limits_{k\to\infty}\|f_{k}\|_{X}$.
Another property that will be important for us is that of the absolutely continuous norm. We follow again the formulation in [25].
5.4 Definition.
Let $X$ be a Banach function space over a $\sigma$-finite measure space $(\Omega,\mu)$. We say that $X$ has a absolutely continuous norm, if for all $f\in X$ and for all decreasing sequences $(E_{k})_{k=1}^{\infty}$ of measurable sets with $\mu\left(\displaystyle\bigcap_{k=1}^{\infty}E_{k}\right)=0$, there holds $\lim\limits_{k\to\infty}\|f\chi_{E_{k}}\|_{X}=0$.
Let us now remark that the weighted variable Lebesgue spaces we consider in this paper are Banach functions spaces satisfying the additional properties.
5.5 Lemma.
Let $p(\cdot)\in\mathscr{P}$ and let $w$ be a weight on $\mathbb{R}^{n}$. Then, $L^{p(\cdot)}(\mathbb{R}^{n})$ is a Banach function space over $\mathbb{R}^{n}$ (equipped with Lebesgue measure) that is reflexive as a Banach space with Banach dual $L^{p^{\prime}(\cdot)}(w^{-1})$, satisfies the Fatou property and has absolutely continuous norm.
Proof.
First of all, it is well-known that $L^{p(\cdot)}(\mathbb{R}^{n})$ is a Banach function space over $\mathbb{R}^{n}$ that satisfies the Fatou property, see for instance [13, Lemma 3.2.8 (b)], and has absolutely continuous norm, see for example [12, p. 73]. Then, by [25, Proposition 3.17] we immediately obtain that $L^{p(\cdot)}(w)$ is a Banach function space over $\Omega$ that has the Fatou property. Furthermore, by combining [25, Proposition 3.12] with [25, Proposition 3.17] we deduce that $L^{p(\cdot)}(w)$ has absolutely continuous norm. Lastly, it is well-known that the Köthe dual of $L^{p(\cdot)}(\mathbb{R}^{n})$ (see [25] for the precise, general definition of the Köthe dual) is
$(L^{p(\cdot)}(\mathbb{R}^{n}))^{\prime}=L^{p^{\prime}(\cdot)}(\mathbb{R}^{n}),$
see for instance [12, p. 73]. Thus, by [25, Proposition 3.17] we obtain that the Köthe dual of $L^{p(\cdot)}(w)$ is
$(L^{p(\cdot)}(w))^{\prime}=L^{p^{\prime}(\cdot)}(w^{-1}).$
In particular, $(L^{p(\cdot)}(w))^{\prime}$ has also absolutely continuous norm. Combining [25, Proposition 3.12] with [25, Corollary 3.16] we immediately deduce that $L^{p(\cdot)}(w)$ is reflexive. From [25, Proposition 3.15 (a)] we finally obtain that the Banach dual of $L^{p(\cdot)}(w)$ is $L^{p^{\prime}(\cdot)}(w^{-1})$. ∎
As proved in [25, Proposition 2.5], the saturation property has several equivalent reformulations. One of them is so important for us that we state it explicitly:
5.6 Lemma.
Let $X$ be a Banach function space over a $\sigma$-finite measure space $(\Omega,\mu)$. Then, there exists an increasing sequence of measurable sets $(A_{k})^{\infty}_{k=1}$ such that $\mu(A_{k})<\infty$ and $\chi_{A_{k}}\in X$ for all $k=1,2,\ldots$, and $\Omega=\bigcup_{k=1}^{\infty}A_{k}$.
Proof.
By [25, Proposition 2.5 (ii)] we have that there exists an increasing sequence of measurable sets $(F_{k})^{\infty}_{k=1}$ such that $\chi_{F_{k}}\in X$ for all $k=1,2,\ldots$, and $\Omega=\bigcup_{k=1}^{\infty}F_{k}$. Moreover, since $\Omega$ is $\sigma$-finite, we can write $\Omega=\bigcup_{k=1}^{\infty}E_{k}$, where $(E_{k})^{\infty}_{k=1}$ is an increasing sequence of measurable sets with $\mu(E_{k})<\infty$, for all $k=1,2,\ldots$. Set $A_{k}:=F_{k}\cap E_{k}$, $k=1,2,\ldots$. Then, $(A_{k})^{\infty}_{k=1}$ is an increasing sequence of measurable sets with $\mu(A_{k})<\infty$, for all $k=1,2,\ldots$ and $\Omega=\bigcup_{k=1}^{\infty}A_{k}$. Moreover, since $0\leq\chi_{A_{k}}\leq\chi_{F_{k}}$, the ideal property yields $\chi_{A_{k}}\in X$, for all $k=1,2,\ldots$. ∎
We now turn to an important embedding property of weighted variable Lebesgue spaces, whose underlying weight satisfies an additional assumption present in our context.
5.7 Definition.
Let $(\Omega,\mu)$ be a $\sigma$-finite measure space. Following [20], we say that a sequence $(f_{k})^{\infty}_{k=1}$ of measurable functions on $\Omega$ converges locally in measure to a measurable function $f$ on $\Omega$, if for any measurable set $E\subseteq\Omega$ with $\mu(E)<\infty$ we have $f_{k}|_{E}\to f|_{E}$ as $k\to\infty$ in measure.
The local convergence in measure was characterized through a metric on the space of measurable functions in [20, Proposition B.6].
5.8 Lemma.
Let $(\Omega,\mu)$ be a $\sigma$-finite measure space. Let $(A_{k})^{\infty}_{k=1}$ be any sequence of pairwise disjoint measurable sets of finite positive measure covering $\Omega$. Consider the map $d:L^{0}(\Omega)\times L^{0}(\Omega)\to[0,\infty)$ given by
$d(f,g):=\sum_{k=1}^{\infty}\frac{1}{2^{k}|A_{k}|}\int_{A_{k}}\frac{|f-g|}{1+|f-g|}\,\mathrm{d}x.$
By [20, Proposition B.6] we have that $d$ is a well-defined metric on $L^{0}(\Omega)$ making it into a topological vector space. Let $(f_{k})^{\infty}_{k=1}$ be a sequence of measurable functions on $\Omega$ and let $f\in L^{0}(\Omega)$. Then, the following are equivalent:
  1. (a)
    $f_{k}\overset{d}{\to}f$ as $k\to\infty$.
  2. (b)
    $f_{k}\to f$ locally in measure as $k\to\infty$.
  3. (c)
    For any $k=1,2,\ldots$ we have $f_{m}|_{A_{k}}\to f|_{A_{k}}$ as $m\to\infty$ in measure.
Proof.
That (a) implies (b) is just direction $\Rightarrow$ in [20, Proposition B.6 (iii)]. It is clear that (b) implies (c). Finally, the proof of direction $\Leftarrow$ in [20, Proposition B.6 (iii)] shows that (c) implies (a), since in that proof one only uses that $f_{m}|_{A_{k}}\to f|_{A_{k}}$ as $m\to\infty$ in measure, for all $k=1,2,\ldots$. ∎
The following estimate will turn out to be crucial for the embedding of weighted variable Lebesgue spaces into the space of measurable functions equipped with the topology of local convergence in measure.
5.9 Lemma.
Let $p(\cdot)\in\mathscr{P}$ and let $w$ be a weight on $\mathbb{R}^{n}$. Then, there exists a sequence $(A_{k})^{\infty}_{k=1}$ of pairwise disjoint measurable sets of finite measure covering $\mathbb{R}^{n}$, such for all $k=1,2,\ldots$, there exists a constant $C_{k}\in(0,\infty)$ with
$\int_{A_{k}}|f(x)|\,\mathrm{d}x\leq C_{k}\|f\|_{L^{p(\cdot)}(w)},\quad\forall f\in L^{p(\cdot)}(w).$
Proof.
By Lemma 5.6, applied for $L^{p^{\prime}(\cdot)}(w^{-1})$, there exists an increasing sequence of measurable sets $(B_{k})_{k=1}^{\infty}$ such that $\mu(B_{k})<\infty$ and $\chi_{B_{k}}\in L^{p^{\prime}(\cdot)}(w^{-1})$ covering $\mathbb{R}^{n}$. Set $A_{1}:=B_{1}$ and
$A_{k}:=B_{k}\setminus\left(\bigcup_{m=1}^{k-1}B_{m}\right),\quad k=2,3,\ldots.$
Then, $A_{1},A_{2},\ldots$ are measurable, pairwise disjoint, have finite measure and cover $\mathbb{R}^{n}$. Furthermore, $\chi_{A_{k}}\in L^{p^{\prime}(\cdot)}(w^{-1})$ for all $k=1,2,\ldots$ by the ideal property.
Fix now $k\geq 1$. Let $f\in L^{p(\cdot)}(w)$ be arbitrary. By Hölder’s inequality in the variable exponent setting, Lemma 3.2, for $p(\cdot)$ and $p^{\prime}(\cdot)$ we have
$\int_{A_{k}}|f(x)|\,\mathrm{d}x$ $=\int_{\mathbb{R}^{n}}|f(x)w(x)|\chi_{A_{k}}(x)w^{-1}(x)\,\mathrm{d}x\leq C\|fw\|_{L^{p(\cdot)}(\mathbb{R}^{n})}\|\chi_{A_{k}}w^{-1}\|_{L^{p^{\prime}(\cdot)}(\mathbb{R}^{n})}$
$=C_{k}\|f\|_{L^{p(\cdot)}(w)},$
where $C_{k}:=C\|\chi_{A_{k}}\|_{L^{p^{\prime}(\cdot)}(w^{-1})}$ is finite. ∎
Now we can state the embedding.
5.10 Proposition.
Let $p(\cdot)\in\mathscr{P}$ and let $w$ be a weight on $\mathbb{R}^{n}$. Then, the inclusion $L^{p(\cdot)}(w)\subseteq L^{0}(\mathbb{R}^{n})$ is a continuous embedding.
Proof.
This is proved identically to [20, Proposition 3.3.1.3], by using Lemma 5.8 (c) in tandem with Lemma 5.9, instead of assumption [20, Definition 3.3.1.1 (iv)]. ∎
5.B. The Cobos–Fernández-Cabrera–Martínez theorem
5.11 Definition.
Let $\bar{A}=(A_{0},A_{1})$ be a pair of Banach spaces. We say that $\bar{A}$ is a Banach couple, if there is some Hausdorff topological vector space $Y$ such that both $A_{0}$ and $A_{1}$ are continuously embedded in $Y$.
We now recall the result of Cobos–Fernández-Cabrera–Martínez [7] about complex interpolation for bilinear compact operators. To state it, we need two additional pieces of notations: if $\bar{A}=(A_{0},A_{1})$ is a Banach couple, the we denote by $A_{j}^{\circ}$ the norm topological closure of $A_{0}\cap A_{1}$ in $A_{j}$, for all $j=0,1$. Moreover, we denote by ${\bf\mathcal{B}}(\bar{A}\times\bar{B},\bar{E})$ the family of bilinear operators $T$ defined on $(A_{0}\cap A_{1})\times(B_{0}\cap B_{1})$ with values in $E_{0}\cap E_{1}$ that satisfy the following
$\|T(a,b)\|_{E_{j}}\leq M_{j}\|a\|_{A_{j}}\|b\|_{B_{j}},\quad a\in A_{0}\cap A_{1},\quad b\in B_{0}\cap B_{1},\quad j=0,1,$
where $M_{j}$ are positive constants.
5.12 Theorem ([7], Theorem 3.2).
Let $\bar{A}=(A_{0},A_{1})$, $\bar{B}=(B_{0},B_{1})$ be Banach couples. Let $\bar{E}=(E_{0},E_{1})$ be a couple of Banach function spaces on a $\sigma$-finite measure space $(\Omega,\mu)$ having the Fatou property. Let $T\in{\bf\mathcal{B}}(\bar{A}\times\bar{B},\bar{E})$. If $T:A_{1}^{\circ}\times B_{1}^{\circ}\to E_{1}$ compactly and $E_{1}$ has absolutely continuous norm, then for any $0<\theta<1$, $T$ may be uniquely extended to a compact bilinear operator from $A_{0},A_{1}]_{\theta}\times[B_{0},B_{1}]_{\theta}$ to $E_{0},E_{1}]_{\theta}$.
Several remarks are in order.
5.13 Remark.
In the original statement of [7, Theorem 3.2], the compactness property was assumed for $T:A_{0}^{\circ}\times B_{0}^{\circ}\to E_{0}$. However, because of the symmetry of the interpolation operator (see for example [3, Theorem 4.2.1]), this is equivalent to the formulation given here.
The definition of Banach function spaces in [7] includes more assumptions than our own. However, a careful examination of the very detailed exposition of the proof of [7, Theorem 3.2] in [20] (and the references therein) shows that the definition of Banach function spaces assumed here is sufficient.
In particular, let us remark that whenever one needs a sequence $(A_{k})^{\infty}_{k=1}$ of measurable sets having finite measure such that $\chi_{A_{k}}$ lies in the considered Banach space, for al $k=1,2,\ldots$ and $\bigcup_{k=1}^{\infty}A_{k}=\Omega$, such as in the proof of [20, Proposition 3.3.3.5], one can use Lemma 5.6. That is, one does not need to assume that the characteristic function of any measurable set of finite measure lies in the Banach space.
5.14 Remark.
For $j=0,1$, let $p_{j}(\cdot)\in\mathscr{P}$ and $w_{j}$ be weights on $\mathbb{R}^{n}$, $j=0,1$. Then, by Proposition 5.10 we have that $(L^{p_{0}(\cdot)}(w_{0}),L^{p_{1}(\cdot)}(w_{1}))$ is a Banach couple.
6. Proof of abstract main result
In this section we finalize the proof of our abstract main result, Theorem 1.1.
The last ingredient we need is the following interpolation result which can be found in [30] (see also [15, Theorem 3.2]).
6.1 Theorem ([30], Theorems 3.4.1 and 3.4.2).
Let $0<\theta<1$. For $j=0,1$, let $p_{j}(\cdot)\in\mathscr{P}$ and $w_{j}$ be weights on $\mathbb{R}^{n}$, $j=0,1$. Then
$L^{p_{0}(\cdot)}(w_{0}),L^{p_{1}(\cdot)}(w_{1})]_{\theta}=L^{p_{\theta}(\cdot)}(w_{\theta}),$
where
$\frac{1}{p_{\theta}(\cdot)}=\frac{1-\theta}{p_{0}(\cdot)}+\frac{\theta}{p_{1}(\cdot)}\quad\text{and}\quad w_{\theta}=w_{0}^{1-\theta}w_{1}^{\theta}.$
Moreover, the norms of the spaces $L^{p_{0}(\cdot)}(w_{0}),L^{p_{1}(\cdot)}(w_{1})]_{\theta}$ and $L^{p_{\theta}(\cdot)}(w_{\theta})$ are equivalent.
We make some remarks about Theorem 6.1:
6.2 Remark.
From Remark 5.14 we know that $(L^{p_{0}(\cdot)}(w_{0}),L^{p_{1}(\cdot)}(w_{1}))$ is a Banach couple.
The statement of [30, Theorem 3.4.1] includes also the assumptions that $w_{j}\in L_{\text{loc}}^{p_{j}(\cdot)}$ and $w_{j}^{-1}\in L_{\text{loc}}^{p_{j}^{\prime}(\cdot)}$. However, in view of Lemma 5.5, a careful examination of the proof given there shows that this assumption is superfluous.
Also, the statement of [30, Theorem 3.4.2] includes the assumptions that $w_{j}\in L_{\text{loc}}^{p_{j}(\cdot)}$ and $w_{j}^{-1}\in L_{\text{loc}}^{p_{j}^{\prime}(\cdot)}$. However, as commented in the proof of [15, Theorem 3.2], the only property used in the proof is that weighted variable Lebesgue spaces are Banach lattices, which in view of Lemma 5.5 does not require these extra assumptions.
We have now all ingredients to prove Theorem 1.1.
Proof of Theorem 1.1.
We prove the theorem in the case $t\in(1,\infty)$. The proof of the remaining case $t=1$ follows by similar arguments with the main difference that below one applies Lemma 4.1 instead of Lemma 4.2.
Let $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, $q(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ such that
$t<\min\{(p_{j})_{-}:~1,2\}$
and
$\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma.$
Let also $\vec{w}$ be a 2-tuple of weights with $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$. We will show that $T:L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})\to L^{q(\cdot)}(\nu_{\vec{w}})$ compactly.
To see that, we first apply Lemma 4.2: It delivers $\theta>0$, $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, $q_{0}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$, such that
$t<\min\{(p_{0},j)_{-}:~j=1,2\},$
$\frac{1}{\vec{p}(\cdot)}=\frac{1-\theta}{\vec{p}_{0}(\cdot)}+\frac{\theta}{\vec{p}_{1}(\cdot)},\quad\frac{1}{q(\cdot)}=\frac{1-\theta}{q_{0}(\cdot)}+\frac{\theta}{q_{1}(\cdot)},\quad\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$
and a 2-tuple of weights $\vec{w}_{0}$ with $\vec{w}=\vec{w}_{0}^{1-\theta}\vec{w}_{1}^{\theta}$, such that $\vec{w}_{0}\in\mathcal{A}^{\gamma}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}$ and $\vec{w}_{0}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}_{0}(\cdot)}{t},\frac{q_{0}(\cdot)}{t}}$.
Set now
$A_{j}:=L^{p_{j,1}(\cdot)}(w_{j,1}),\quad B_{j}:=L^{p_{j,2}(\cdot)}(w_{j,2}),\quad E_{j}:=L^{q_{j}(\cdot)}(\nu_{\vec{w}_{j}}),\quad j=0,1.$
By Lemma 5.5 and Remark 5.14 we have that $A_{j},B_{j},E_{j}$, $j=0,1$ satisfy all assumptions of Theorem 5.12. Moreover, since by assumption the operator $T:A_{1}\times B_{1}\to E_{1}$ is compact and $A_{1}^{\circ}$, respectively $B_{1}^{\circ}$, is a subspace of $A_{1}$, respectively $B_{1}$, we easily deduce from the equivalent characterizations of compactness of bilinear operators (see for instance [2, Proposition 1]) that also the restriction $T:A_{1}^{\circ}\times B_{1}^{\circ}\to E_{1}$ is compact. Therefore, Theorem 5.12 yields
$T:[A_{0},A_{1}]_{\theta}\times[B_{0},B_{1}]_{\theta}\to[E_{0},E_{1}]_{\theta}\quad\text{compactly}.$
Applying Theorem 6.1 we finally deduce
$T:L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})\to L^{q(\cdot)}(\nu_{\vec{w}})\quad\text{compactly}.$
7. Applications
In this section, we provide several applications of our compact extrapolation Theorem 1.1 for bilinear operators. In particular, we will show the compactness of commutators of bilinear $\omega$-Calderón–Zygmund operators and bilinear fractional integral operators.
7.A. Notation for multilinear commutators
Given some $m$-linear operator $T$ accepting as arguments $m$-tuples $\vec{f}=(f_{1},\ldots,f_{m})$ of appropriate functions on $\mathbb{R}^{n}$ and some appropriately smooth function $b$ on $\mathbb{R}^{n}$, we define for each $j=1,\ldots,m$ the commutator
$T,b]_{e_{j}}(\vec{f})(x):=b(x)T(\vec{f})(x)-T(f_{1},\ldots,bf_{j},\ldots,f_{m})(x).$
As we will see below, in order to apply our extrapolation Theorem 1.1 we require that the pointwise multiplier $b$ belongs to the spaces
$\operatorname{BMO}(\mathbb{R}^{n}):=\Big\{f:\mathbb{R}^{n}\to\mathbb{C}\ \Big|\ \|f\|_{\operatorname{BMO}(\mathbb{R}^{n})}:=\sup_{Q\subset\mathbb{R}^{n}}\langle|f-\langle f\rangle_{Q}|\rangle_{Q}<\infty\Big\}$
or
$\operatorname{CMO}(\mathbb{R}^{n}):=\overline{C_{c}^{\infty}(\mathbb{R}^{n})}^{\operatorname{BMO}(\mathbb{R}^{n})}.$
Moreover, given a $m$-tuple $\vec{b}=(b_{1},\ldots,b_{m})$ of appropriately smooth functions on $\mathbb{R}^{n}$, we consider the sum commutator
$T,\vec{b}]_{\Sigma}(\vec{f}):=\sum_{j=1}^{m}[T,b_{j}]_{e_{j}}(\vec{f}).$
7.B. Notation for maximal operators
For $\delta,\varepsilon,\alpha>0$ and $\vec{r}=(r_{1},\ldots,r_{m})\in(1,\infty)^{m}$ we consider the maximal operators
$M^{\#}_{\delta}(f)(x):=\sup_{\begin{subarray}{c}Q\text{ cube in }\mathbb{R}^{n}\\ x\in Q\end{subarray}}\left(\frac{1}{|Q|}\int_{Q}||f(y)|^{\delta}-\langle|f|^{\delta}\rangle_{Q}|\,\mathrm{d}y\right)^{1/\delta},$
$M_{\varepsilon}(f)(x):=\sup_{\begin{subarray}{c}Q\text{ cube in }\mathbb{R}^{n}\\ x\in Q\end{subarray}}\left(\frac{1}{|Q|}\int_{Q}|f(y)|^{\varepsilon}\,\mathrm{d}y\right)^{1/\varepsilon},$
$\mathcal{M}_{\alpha}(\vec{f})(x):=\sup_{\begin{subarray}{c}Q\text{ cube in }\mathbb{R}^{n}\\ x\in Q\end{subarray}}\frac{1}{|Q|^{m-\frac{\alpha}{n}}}\prod_{j=1}^{m}\int_{Q}|f_{j}(y_{j})|\mathrm{d}y_{j},$
$\mathcal{M}(\vec{f})(x):=\sup_{\begin{subarray}{c}Q\text{ cube in }\mathbb{R}^{n}\\ x\in Q\end{subarray}}\prod_{j=1}^{m}\frac{1}{|Q|}\int_{Q}|f_{j}(y_{j})|\mathrm{d}y_{j},$
and
$\mathcal{M}_{\alpha,\mathrm{L(logL)}}(\vec{f})(x):=\sup_{\begin{subarray}{c}Q\text{ cube in }\mathbb{R}^{n}\\ x\in Q\end{subarray}}|Q|^{\alpha/n}\prod_{k=1}^{m}\|f_{k}\|_{\mathrm{L(logL),Q}},$
where $\|f\|_{\mathrm{L(logL),Q}}$ is the usual $\mathrm{L(\log L)}$ Luxemburg norm of $f\mathbf{1}_{Q}$ over $Q$ with respect to normalized Lebesgue measure on $Q$, concretely
$\|f\|_{\mathrm{L(logL),Q}}:=\inf\left\{\lambda>0:~\frac{1}{|Q|}\int_{Q}\frac{1}{\lambda}\log\left(e+\frac{|f(y)|}{\lambda}\,\mathrm{d}y\right)\leq 1\right\}.$
7.C. Bilinear $\omega$-Calderón–Zygmund operators
Let $\omega:[0,\infty)\rightarrow[0,\infty)$ be a modulus of continuity, that is, $\omega$ is increasing, subadditive and $\omega(0)=0$. A function $K(x,y_{1},\ldots,y_{m})$ on $(\mathbb{R}^{n})^{{m+1}}$ defined away from the diagonal $x=y_{1}=\ldots=y_{m}$ is said to be an $\omega$-Calderón–Zygmund kernel if the following size and smoothness conditions are satisfied:
$|K(x,y_{1},\ldots,y_{m})|$ $\lesssim\frac{1}{(\sum_{j=1}^{m}|x-y_{j}|)^{mn}},$
$|K(x,y_{1},\ldots,y_{m})-K(x^{\prime},y_{1},\ldots,y_{m})|$ $\lesssim\frac{1}{(\sum_{j=1}^{m}|x-y_{j}|)^{mn}}\omega\bigg(\frac{|x-x^{\prime}|}{\sum_{j=1}^{m}|x-y_{j}|}\bigg),$
whenever $|x-x^{\prime}|\leq\frac{1}{2}\max_{j=1,\ldots,m}|x-y_{j}|$, and for each $i=1,\dots,m,$
$|K(x,y_{1},\ldots,y_{i},\ldots,y_{m})-K(x,y_{1},\ldots,y_{i}^{\prime},\ldots,y_{m})|\lesssim\frac{1}{(\sum_{j=1}^{m}|x-y_{j}|)^{mn}}\omega\bigg(\frac{|y_{i}-y_{i}^{\prime}|}{\sum_{j=1}^{m}|x-y_{j}|}\bigg),$
whenever $|y_{i}-y_{i}^{\prime}|\leq\frac{1}{2}\max_{j=1,\ldots,m}|x-y_{j}|$.
The $m$-linear operator $T:\mathcal{S}(\mathbb{R}^{n})\times\cdots\times\mathcal{S}(\mathbb{R}^{n})\rightarrow\mathcal{S}^{\prime}(\mathbb{R}^{n})$ is said to be an $\omega$-Calderón–Zygmund operator if it extends to a bounded multilinear operator from $L^{q_{1}}(\mathbb{R}^{n})\times\cdots\times L^{q_{m}}(\mathbb{R}^{n})$ to $L^{q}(\mathbb{R}^{n})$ for some $\frac{1}{q}=\frac{1}{q_{1}}+\cdots+\frac{1}{q_{m}}$ with $1<q_{1},\dots,q_{m}<\infty$, and there exists an $\omega$-Calderón–Zygmund kernel $K$ such that
$T(\vec{f})(x):=\int_{(\mathbb{R}^{n})^{m}}K(x,y_{1},\ldots,y_{m})f(y_{1})\ldots f(y_{m})\,\mathrm{d}y_{1}\ldots\mathrm{d}y_{m},$
for all $x\notin\cap_{j=1}^{m}\operatorname{supp}f_{j}$, $j=1,\dots,m$.
We recall that a modulus of continuity $\omega$ is said to satisfy the Dini condition, denoted by $\omega\in$ Dini, if
$\|\omega\|_{\text{Dini}}:=\int_{0}^{1}\omega(t)\frac{\mathrm{d}t}{t}<\infty.$
Notice that the modulus of continuity $\omega(t)=t^{\varepsilon}$ with $\varepsilon>0$ satisfies the Dini condition. This example appeared in the work of Grafakos and Torres [16], where they studied the (standard) Calderón–Zygmund operator. It is worth to mention that for the general $\omega$, the linear $\omega$-Calderón–Zygmund operator was studied by Yabuta in [35]. A further extension of this result was obtained by Maldonado and Naibo [28] in the bilinear case.
The authors of [5, Theorem 5.1] proved the compactness of commutators of m-linear $\omega$-Calderón–Zygmund operators on weighted Lebesgue spaces with constant exponents. We point out that one part of the proof of [5, Theorem 5.1] relied on showing the following unweighted compactness result.
7.1 Theorem ([5]).
Let $b\in\mathrm{CMO}(\mathbb{R}^{n})$, $1<p_{1},\ldots,p_{m}<\infty$ and $0<p<\infty$ with $\frac{1}{p}:=\sum_{j=1}^{m}\frac{1}{p_{j}}$. Then, $T,b]_{e_{j}}$ maps $L^{p_{1}}(\mathbb{R}^{n})\times\ldots\times L^{p_{m}}(\mathbb{R}^{n})$ into $L^{p}(\mathbb{R}^{n})$ compactly for every $j=1,\ldots,m$.
We will extend Theorem 7.1 to the weighted variable Lebesgue spaces in the bilinear setting. To this end, we need the following two weighted boundedness results.
7.2 Theorem.
If $T$ is a bilinear $\omega$-Calderón–Zygmund operator, then for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ with $\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}$ and for all pairs of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot)}$, it holds that $T$ maps $L^{p_{1}(\cdot)}(w_{1})\times\ldots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{p(\cdot)}(\nu_{\vec{w}})$ boundedly.
Proof.
The proof is almost the same to the one given in [33, Theorem 2.8]. In particular, the main difference is that instead of the pointwise estimate of [33, Proposition 7.2] we make use of the following one (see [27, Theorem 6.1])
$M^{\#}_{\delta}(T(\vec{f}))(x)\lesssim\mathcal{M}(\vec{f})(x),$
where $0<\delta<1/m$. ∎
7.3 Remark.
We notice that one of the main ingredients of the proof of [33, Theorem 2.8] was the characterization of $\mathcal{A}_{\vec{p}(\cdot)}$ weights by the bilinear maximal operator $\mathcal{M}$ (see [33, Theorem 2.4]). According to [33, Remark 2.7], one can obtain the multilinear version of this latter result. We also refer to the work [6, Theorem 2.2], in which the characterization of the $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ weights was established with the aid of the multilinear fractional maximal operator $\mathcal{M}_{\alpha}$.
7.4 Theorem.
Let $b\in\mathrm{BMO}(\mathbb{R}^{n})$. Let $t$ be a fixed constant with $t>1$. Let $j\in\{1,\ldots,m\}$. Then, for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ with
$\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}$
and
$t<\min\{(p_{j})_{-}:~j=1,\ldots,m\},$
and for $m$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}_{\frac{\vec{p}(\cdot)}{t}}$, it holds that $T,b]_{e_{j}}$ maps $L^{p_{1}(\cdot)}(w_{1})\times\ldots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{p(\cdot)}(\nu_{\vec{w}})$ boundedly.
Proof.
The proof is similar to [6, Theorem 2.4]. Here, we provide the necessary changes. In particular, let $0<\delta<\varepsilon<1/m$. Then, by [27, page 95 and Theorem 7.1] we have the pointwise estimate
$M^{\#}_{\delta}([T,b]_{e_{j}}(\vec{f}))(x)$ $\lesssim\|b\|_{\mathrm{BMO}(\mathbb{R}^{n})}(M_{\varepsilon}(T(\vec{f}))(x)+\mathcal{M}_{\mathrm{L(log(L))}}(\vec{f})(x))$
$\lesssim\|b\|_{\mathrm{BMO}(\mathbb{R}^{n})}(M_{\frac{1}{m}}(T(\vec{f}))(x)+\mathcal{M}_{\mathrm{L(log(L))}}(\vec{f})(x)),$
where the last inequality follows by using Jensen’s inequality. The rest of the proof proceeds in a similar fashion as in [6, Theorem 2.4] with the difference that instead of [6, Theorem 2.3] we apply Theorem 7.2; we skip the details. ∎
Now, we are in position to apply our abstract main result Theorem 1.1 which gives immediately the following.
7.5 Theorem.
Let $\vec{b}=(b_{1},b_{2})$ be a pair of functions in $\mathrm{CMO}(\mathbb{R}^{n})$ and $t$ be a fixed constant with $t>1$. Then, for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$ with
$\frac{1}{p(\cdot)}=\frac{1}{p_{1}(\cdot)}+\frac{1}{p_{2}(\cdot)}$
and
$t<\min\{(p_{j})_{-}:~j=1,2\},$
and for all pairs of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}_{\frac{\vec{p}(\cdot)}{t}}$, it holds that $T,b_{1}]_{e_{1}}$, $T,b_{2}]_{e_{2}}$ and $T,\vec{b}]_{\Sigma}$ map $L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})$ into $L^{p(\cdot)}(\nu_{\vec{w}})$ compactly.
Proof.
The first step is to apply Theorem 1.1 for the operators $T,b_{1}]_{e_{1}}$ and $T,b_{2}]_{e_{2}}$, noticing that $\gamma=0$.
Part (1) of the assumptions of Theorem 1.1 for $T,b_{1}]_{e_{1}}$ and $T,b_{2}]_{e_{2}}$ follows from the conclusion of Theorem 7.4 in the bilinear case (with an obvious change of notation).
Choose now constants $0<x_{1},x_{2}<1$ such that
$0<x_{1}+x_{2}<1.$
Then, Theorem 7.1 implies that part (2) of the assumptions of Theorem 1.1 is satisfied with the particular choices $\vec{p}_{1}(\cdot):=\left(\frac{1}{x_{1}},\frac{1}{x_{2}}\right)$, $p_{1}(\cdot):=\frac{1}{x_{1}+x_{2}}$ and $\vec{w}_{1}:=(1,1)$ for each of the operators $T,b_{1}]_{e_{1}}$ and $T,b_{2}]_{e_{2}}$.
Hence, by Theorem 1.1 we deduce the desired conclusion for each of the operators $T,b_{1}]_{e_{1}}$ and $T,b_{2}]_{e_{2}}$. This implies immediately the desired conclusion also for $T,\vec{b}]_{\Sigma}$, since the sum of bilinear compact operators remains compact due to [2, Proposition 2]. ∎
We close this subsection by mentioning that there are several examples of the bilinear $\omega$-Calderón–Zygmund operators (see [5, Subsection 5.1]) which satisfy the assumptions and conclusion of Theorem 7.5.
7.D. Bilinear fractional Calderón–Zygmund operators
Let $m\geq 2$ and $\alpha\in(0,mn)$. A function $K_{\alpha}(x,y_{1},\ldots,y_{m})$ on $(\mathbb{R}^{n})^{m+1}$ defined away from the diagonal $x=y_{1}=\ldots=y_{m}$ is said to be a $m$-linear $\alpha$-fractional Calderón–Zygmund kernel if there exist constants $A,B>0$, such that the following size estimate is satisfied:
$|K_{\alpha}(x,y_{1},\ldots,y_{m})|\leq\frac{A}{(|x-y_{1}|+\ldots+|x-y_{m}|)^{mn-\alpha}},$
as well as the following smoothness estimates hold:
  • It holds
    $|K_{\alpha}(x,y_{1},\ldots,y_{m})-K_{\alpha}(x^{\prime},y_{1},\ldots,y_{m})|\leq\frac{B|x-x^{\prime}|}{(|x-y_{1}|+\ldots+|x-y_{m}|)^{mn-\alpha+1}},$
    whenever $|x-x^{\prime}|\leq\frac{1}{2}\max_{j=1,\ldots,m}|x-y_{j}|$.
  • For each $j=1,\ldots,m$, one has
    $|K_{\alpha}(x,y_{1},\ldots,y_{j},\ldots,y_{m})-K_{\alpha}(x,y_{1},\ldots,y_{j}^{\prime},\ldots,y_{m})|\leq\frac{B|y_{j}-y_{j}^{\prime}|}{(|x-y_{1}|+\ldots+|x-y_{m}|)^{mn-\alpha+1}},$
    whenever $|y_{j}-y_{j}^{\prime}|\leq\frac{1}{2}\max_{k=1,\ldots,m}|x-y_{k}|$.
As noted in [1] in the special case $m=2$, given such a kernel $K_{\alpha}$, the $m$-linear operator $T_{\alpha}$ defined by
$T_{\alpha}(\vec{f})(x):=\int_{(\mathbb{R}^{n})^{m}}K_{\alpha}(x,y_{1},\ldots,y_{m})f_{1}(y_{1})\ldots f_{m}(y_{m})\,\mathrm{d}y_{1}\ldots\mathrm{d}y_{m},\quad x\in\mathbb{R}^{n}$
is well defined whenever $f_{j}$, $j=1,\ldots,m$ are for example bounded functions with compact support on $\mathbb{R}^{n}$.
As observed in [1, page 3] in the special case $m=2$ (see also [5, Theorem 5.8]), an example of such an operator is the $m$-linear fractional integral operator $\mathcal{I}_{\alpha}$, also known as the $m$-linear Riesz potential, associated to the kernel
$K_{\alpha}(x,y_{1},\ldots,y_{m}):=\frac{1}{(|x-y_{1}|+\ldots+|x-y_{m}|)^{mn-\alpha}}.$
In fact, for every such operator $T_{\alpha}$ one has by the size estimate of the kernel
(7.6) $|T_{\alpha}(\vec{f})(x)|\leq A\,\mathcal{I}_{\alpha}(|f_{1}|,\ldots,|f_{m}|)(x).$
It was shown in [19] that $\mathcal{I}_{\alpha}$ extends to a bounded $m$-linear operator $L^{p_{1}}(\mathbb{R}^{n})\times\ldots\times L^{p_{m}}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})$ whenever $1<p_{1},\ldots,p_{m}<\infty$, $0<q<\infty$ and $\frac{\alpha}{n}=\frac{1}{p}-\frac{1}{q}$, where $\frac{1}{p}:=\sum_{j=1}^{m}\frac{1}{p_{j}}$. Thus, the same holds for arbitrary $T_{\alpha}$.
A result on the compactness of commutators of $T_{\alpha}$ on unweighted Lebesgue spaces in the bilinear case was first established in [1, Theorem 2.1]. Subsequently, [5, Theorem 5.8] extended [1, Theorem 2.1] to the general $m$-linear setting and (constant exponent) weighted Lebesgue spaces and explicitly allowed the target space to be only a quasi-Banach space. Let us note that while [5, Theorem 5.8] is formally stated only for $\mathcal{I}_{\alpha}$, the authors of [5] observe that the statement and their proof hold for arbitrary $T_{\alpha}$. Here we only need the unweighted version of [5, Theorem 5.8].
7.7 Theorem ([1, 5]).
Let $b\in\mathrm{CMO}(\mathbb{R}^{n})$. Let $1<p_{1},\ldots,p_{m}<\infty$ and let $0<q<\infty$ such that $\frac{1}{p}-\frac{1}{q}=\frac{\alpha}{n}$, where $\frac{1}{p}:=\sum_{j=1}^{m}\frac{1}{p_{j}}$. Then, $T_{\alpha},b]_{e_{j}}$ maps $L^{p_{1}}(\mathbb{R}^{n})\times\ldots\times L^{p_{m}}(\mathbb{R}^{n})$ into $L^{q}(\mathbb{R}^{n})$ compactly for every $j=1,\ldots,m$.
Our goal consists in generalizing Theorem 7.7 to weighted variable exponent Lebesgue spaces in the bilinear case. To achieve that, we first need a boundedness result over such spaces. For the special case of the $m$-Riesz potential, this is just [6, Theorem 2.4]. Our proof of the general case is only a slight adaptation of [6, Theorem 2.4].
7.8 Theorem.
Let $b\in\mathrm{BMO}(\mathbb{R}^{n})$. Let $t$ be a fixed constant with $t>1$ and $\gamma:=\frac{\alpha}{n}<\frac{m}{t}$. Let $j\in\{1,\ldots,m\}$. Then, for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ and $q(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}$ with
$\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)},\quad\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\frac{\alpha}{n}$
and
$t<\min\{(p_{j})_{-}:~j=1,\ldots,m\},$
and for $m$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$, it holds that $T_{\alpha},b]_{e_{j}}$ maps $L^{p_{1}(\cdot)}(w_{1})\times\ldots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{q(\cdot)}(\nu_{\vec{w}})$ boundedly.
Proof.
The proof is almost identical to [6, Theorem 2.4], so we explain only the necessary changes.
First of all, by [6, Theorem 2.3] we have that the $m$-linear Riesz potential $\mathcal{I}_{\alpha}$ maps $L^{p_{1}(\cdot)}(w_{1})\times\ldots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{q(\cdot)}(\nu_{\vec{w}})$ boundedly. Thus, by (7.6) we have that the same holds for arbitrary $T_{\alpha}$.
Let now $0<\delta<\varepsilon<\frac{n}{mn-\alpha}$. Then, by [36, Lemma 2.24] we have the pointwise estimate
$M^{\#}_{\delta}([T_{\alpha},b]_{e_{j}}(\vec{f}))(x)\leq C\|b\|_{\mathrm{BMO}(\mathbb{R}^{n})}(M_{\varepsilon}(T_{\alpha}(\vec{f}))(x)+\mathcal{M}_{\alpha,\mathrm{L(log(L))}}(\vec{f})(x)),$
where the constant $C$ depends only on $n,m$, the constants of $T_{\alpha}$, $\varepsilon$ and $\delta$. We now pick $\varepsilon:=\frac{1}{m}$, which is clearly a permitted choice. The rest of the proof proceeds exactly as in [6, Theorem 2.4]; we omit the details. ∎
Applying now our abstract main result, Theorem 1.1, we deduce immediately the following.
7.9 Theorem.
Let $\vec{b}=(b_{1},b_{2})$ be a pair of functions in $\mathrm{CMO}(\mathbb{R}^{n})$. Set $\gamma:=\frac{\alpha}{n}$. Let $t$ be a fixed constant with $1<t<\frac{2}{\gamma}$. Then, for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$ and $q(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ with
$\frac{1}{p(\cdot)}=\frac{1}{p_{1}(\cdot)}+\frac{1}{p_{2}(\cdot)},\quad\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\frac{\alpha}{n}$
and
$t<\min\{(p_{j})_{-}:~j=1,2\},$
and for all pairs of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$, it holds that $T_{\alpha},b_{1}]_{e_{1}}$, $T_{\alpha},b_{2}]_{e_{2}}$ and $T_{\alpha},\vec{b}]_{\Sigma}$ map $L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})$ into $L^{q(\cdot)}(\nu_{\vec{w}})$ compactly.
Proof.
We first apply Theorem 1.1 for the operators $T_{\alpha},b_{1}]_{e_{1}}$ and $T_{\alpha},b_{2}]_{e_{2}}$, noting that $\gamma<\frac{2}{t}$.
Part (1) of the assumptions of Theorem 1.1 for $T_{\alpha},b_{1}]_{e_{1}}$ and $T_{\alpha},b_{2}]_{e_{2}}$ is just the conclusion of Theorem 7.8 in the bilinear case (with an obvious change of notation).
Pick now constants $0<\widetilde{x}_{1},\widetilde{x}_{2}<1$ such that
$\frac{\alpha}{n}<\widetilde{x}_{1}+\widetilde{x}_{2}<1+\frac{\alpha}{n}.$
Such a choice is possible, since $\frac{\alpha}{n}<2$. Then, Theorem 7.7 yields that part (2) of the assumptions of Theorem 1.1 is satisfied with the particular choices $\vec{p}_{1}(\cdot):=\left(\frac{1}{\widetilde{x}_{1}},\frac{1}{\widetilde{x}_{2}}\right)$, $q_{1}(\cdot):=\frac{1}{\widetilde{x}_{1}+\widetilde{x}_{2}-\frac{\alpha}{n}}$ and $\vec{w}_{1}:=(1,1)$ for each of the operators $T_{\alpha},b_{1}]_{e_{1}}$ and $T_{\alpha},b_{2}]_{e_{2}}$.
Thus, by Theorem 1.1 we deduce the desired conclusion for each of the operators $T_{\alpha},b_{1}]_{e_{1}}$ and $T_{\alpha},b_{2}]_{e_{2}}$. This yields immediately the desired conclusion also for $T_{\alpha},\vec{b}]_{\Sigma}$, since the sum of bilinear compact operators remains compact per [2, Proposition 2]. ∎